An airplane took 3 hours to fly 1800 miles against a headwind. The return trip with the wind took 2 hours. Find the speed of the plane in still air and the speed of the wind.
Answer by lwsshak3(11628)

An airplane took 3 hours to fly 1800 miles against a headwind. The return trip with the wind took 2 hours. Find the speed of the plane in still air and the speed of the wind.
***
let x=speed of plane in still air
let c=speed of wind
x+c=speed of plane with the wind
x-c=speed of plane against the wind
travel time*speed=distance
..
3(x-c)=1800
2(x+c)=1800
..
3x-3c=1800
2x+2c=1800
..
6x-6c=3600 (mult eq. by 2)
6x+6c=5400 (mult eq. by 3)
add
12x=9000
x=750
2c=1800-2x=300
c=150
speed of plane in still air=750 mph
speed of wind=150 mph